OpenStudy (anonymous):

Diff EQ: xy'' + y' + xy = 0, x_0 = 1 find a power series representation.

myininaya (myininaya):

so I think I remember something like this.. \[y= \sum_{n=0}^{\infty} a_n x^{n} \] then find y' and y'' from this

OpenStudy (anonymous):

is that for x_0 = 0 though?

myininaya (myininaya):

didn't notice that part \[y=\sum_{n=0}^{\infty} a_n (x-1)^n\]

OpenStudy (anonymous):

ok so i get \[x \sum_{n=2}^{\infty}(n-1)na_{n}(x-1)^{n-2} + \sum_{n=1}^{\infty}na_{n})(x-1)^{n-1} + x \sum_{n=0}^{\infty}a_{n}(x-1)^{n}\]

myininaya (myininaya):

ok and we want to write as one sum n-2=0 let k=n-2 then.. \[x \sum_{n-2=0}^{\infty} (n-1)n a_n(x-1)^{n-2} \\ =x \sum_{k=0}^{\infty} (k+2-1)(k+2)a_{k+2}(x-1)^{k} \\ =\sum_{k=0}^{\infty} (k+1)(k+2)a_{k+2}(x-1)^k x\] then we can do something similar for the middle term n-1=0 let k=n-1 ...

OpenStudy (anonymous):

Ok let me do that real quick and ill post what i get

myininaya (myininaya):

wait...

myininaya (myininaya):

no actually that might work

myininaya (myininaya):

like all of them will have the (x-1)^k thingy

myininaya (myininaya):

but there is still that x in the first and third term ...

OpenStudy (anonymous):

and then you take out terms to get all the summations to start at the same index?

OpenStudy (anonymous):

you cant combine the x and x-1 in any way can you?

myininaya (myininaya):

i'm rethinking a bit... okay so we have that x multiple for the first and third and we have the the whole (x-1) thingy too... I think I want to write x as (x-1)+1

OpenStudy (anonymous):

Question: Is \(x_0=1\) the center of the series, or is it supposed to be an initial condition? The first makes more sense, seeing as \(x=0\) is a singular point.

OpenStudy (anonymous):

its the center of the series

OpenStudy (anonymous):

Ah got it, didn't notice the second/third comments.

myininaya (myininaya):

\[y=\sum_{n=0}^{\infty} a_n (x-1)^n \\ y'=\sum_{n=1}^\infty a_n n(x-1)^{n-1} \\ y'' =\sum_{n=2}^{\infty} a_n (n-1)n(x-1)^{n-2} \\ \text{ ... } \\ xy=[(x-1)+1] \sum_{n=0}^{\infty} a_n(x-1)^n =\sum_{n=0}^{\infty} a_n(x-1)^{n+1}+\sum_{n=0}^{\infty} a_n(x-1)^{n} \\ xy''=[(x-1)+1] \sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-2} = \\\sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-1} +\sum_{n=2}^{\infty}a_n(n-1)n(x-1)^{n-2}\]

myininaya (myininaya):

\[xy''+y'+xy=0 \\ \text{ gives } \sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-1}+\sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-2} \\ +\sum_{n=1}^{\infty} a_n n(x-1)^{n-1}+\sum_{n=0}^{\infty} a_n(x-1)^{n+1} +\sum_{n=0}^{\infty}a_n (x-1)^n=0\] please tell me if you see any type-os in this mess

OpenStudy (anonymous):

so far so good

myininaya (myininaya):

for first term we want n-1=k so we have n=k+1 so if n=2 then k=1 --- for second term we want n-2=k so we have n=k+2 so if n=2 then k=0 --- for third term we want n-1=k so we have n=k+1 so if n=1 then k=0 --- for forth term we want n+1=k so we have n=k-1 so if n=0 then k=1 --- for firth term we want n=k nothing changes just n to k or you can just change all the k's to n's later whatever... \[xy''+y'+xy=0 \\ \text{ gives } \sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-1}+\sum_{n=2}^{\infty} a_n(n-1)n(x-1)^{n-2} \\ +\sum_{n=1}^{\infty} a_n n(x-1)^{n-1}+\sum_{n=0}^{\infty} a_n(x-1)^{n+1} +\sum_{n=0}^{\infty}a_n (x-1)^n=0\] \[\sum_{k=1}^{\infty} a_{k+1}(k)(k+1)(x-1)^{k}+\sum_{k=0}^{\infty} a_{k+2}(k+1)(k+2)(x-1)^{k} \\ +\sum_{k=0}^{\infty} a_{k+1}(k+1)(x-1)^{k}+\sum_{k=1}^{\infty} a_{k-1}(x-1)^{k}+\sum_{k=0}^{\infty} a_k (x-1)^{k}=0\]

myininaya (myininaya):

now we have the same variable part for each one that variable part being the (x-1)^k thingy

myininaya (myininaya):

we are going to need to get the same starting number for each one so we need to expand some of these just a bit

OpenStudy (anonymous):

I got\[2a_{2} + a_{1} + a_{0} + \sum_{n=1}^{\infty}[a_{n+1}n(n+1)+a_{n+2}(n+1)(n+2) + a_{n+1}(n+1)+a_{n-1} + a_{n}](x-1)^{n}\]

OpenStudy (anonymous):

and it got cut off D:

myininaya (myininaya):

\[\sum_{k=0}^{\infty} a_{k+2}(k+1)(k+2)(x-1)^{k} =2a_2+\sum_{k=1}^{\infty}a_{k+2}(k+1)(k+2)(x-1)^{k}\] \[\sum_{k=0}^{\infty} a_{k+1}(k+1)(x-1)^k=a_1+\sum_{k=1}^\infty a_{k+1}(k+1)(x-1)^k \\ \] \[\sum_{k=0}^{\infty} a_k(x-1)^k=a_0+\sum_{k=1}^{\infty} a_k(x-1)^k\] so I see how you got those first three terms

myininaya (myininaya):

let me see if can put it altogether now

myininaya (myininaya):

\[\sum_{k=1}^{\infty} a_{k+1}(k)(k+1)(x-1)^{k}+\sum_{k=0}^{\infty} a_{k+2}(k+1)(k+2)(x-1)^{k} \\ +\sum_{k=0}^{\infty} a_{k+1}(k+1)(x-1)^{k}+\sum_{k=1}^{\infty} a_{k-1}(x-1)^{k}+\sum_{k=0}^{\infty} a_k (x-1)^{k}=0 \\ \\ \] \[2a_2+a_1+a_0+\sum_{k=1}^{\infty}a_{k+1}(k)(k+1)(x-1)^{k}+ \\ +\sum_{k=1}^{\infty} a_{k+2}(k+1)(k+2)(x-1)^{k}+\sum_{k=1}^{\infty} a_{k+1}(k+1)(x-1)^k + \\ +\sum_{k=1}^{\infty} a_{k-1}(x-1)^k+\sum_{k=1}^{\infty}a_k(x-1)^k=0\]

myininaya (myininaya):

\[2a_2+a_1+a_0+ \\ \sum_{k=1}^{\infty} [a_{k+1}k(k+1)+a_{k+2}(k+1)(k+2)+a_{k+1}(k+1)+a_{k-1}+a_k](x-1)^k=0\]

myininaya (myininaya):

I feel like this problem is so intense

myininaya (myininaya):

so many type-os or mistakes could happen

OpenStudy (anonymous):

this problem sucks lol

OpenStudy (anonymous):

For what it's worth, I'm ending up with the same recurrence.

myininaya (myininaya):

so if k=0 then \[2a_2+a_1+a_0=0 \\ \] now the other k integers greater than 0 occur in the series section there... if k=1,2,3,4,5,... then \[a_{k+1}k(k+1)+a_{k+2}(k+1)(k+2)+a_{k+1}(k+1)+a_{k-1}+a_k=0\] that is great to know

myininaya (myininaya):

\[a_{k-1}+a_k+a_{k+1}[k(k+1)+(k+1)]+a_{k+2}(k+1)(k+2)=0 \\ a_{k-1}+a_k+a_{k+1}(k+1)^2+a_{k+2}(k+1)(k+2)=0 \\ a_{k+2}=\frac{-a_{k-1}-a_k-(k+1)^2a_{k+1}}{(k+1)(k+2)} ; k=1,2,3,...\] ... \[a_2=\frac{-a_1-a_0}{2} \] so if k=1... \[a_3=\frac{-a_0-a_1-4a_2}{2(4)}=\frac{\frac{-a_1-a_0}{2}-2a_2}{4}=\frac{a_2-2a_2}{4}=\frac{-a_2}{4}\]

myininaya (myininaya):

oops 1+2 is 3 not 4

OpenStudy (anonymous):

question

myininaya (myininaya):

so should be \[a_3=\frac{-a_2}{3}\]

myininaya (myininaya):

what is the question

OpenStudy (anonymous):

the (k+1)^2 how'd you get that

OpenStudy (anonymous):

nevermind, figured it out. Sorry lol

myininaya (myininaya):

I combined like terms.. \[a_{k+1}k(k+1)+a_{k+1}(k+1)=a_{k+1}(k(k+1)+(k+1)) \\ =a_{k+1}(k+1)(k+1)=a_{k+1}(k+1)^2\] \[a_{k+1}k(k+1)+a_{k+1}(k+1) \\ a_{k+1}(k+1)[k+1] \\ a_{k+1}(k+1)^2 \text{ or you can call it factoring }\]

myininaya (myininaya):

so we found a pretty form for a_3 \[a_3=\frac{-a_2}{3}\] ... let's see what happens when we plug in 2 for k... \[ a_{k+2}=\frac{-a_{k-1}-a_k-(k+1)^2a_{k+1}}{(k+1)(k+2)} ; k=1,2,3,... \\ \] \[a_4=\frac{-a_{1}-a_2-9a_3}{3(5)} ...\] we can write a_3 in temrs of a_2... \[a_4=\frac{-a_1-a_2+3a_2}{3(5)}=\frac{-a_1+2a_2}{3(5)}=\frac{-a_1+2a_2}{15}\]

myininaya (myininaya):

omg I hate addition lol \[a_4=\frac{-a_1+2a_2}{3(4)}=\frac{-a_1+2a_2}{12}\] plug in 3 for k... \[a_5=\frac{-a_2-a_3-16 a_4}{4(5)}\]

myininaya (myininaya):

\[a_5=\frac{-a_2+\frac{a_2}{3}-16 \cdot \frac{-a_1+2a_2}{12}}{20} =\frac{-12a_2+4a_2-16(-a_1+2a_2)}{12(20)} \\ a_5=\frac{ -40a_2+16a_1}{12(20)}=\frac{-10a_2+4a_1}{3(20)}=\frac{-5a_2+2a_1}{3(10)}=\frac{-5a_2+2a_1}{30}\]

myininaya (myininaya):

I'm not seeing a pattern

myininaya (myininaya):

are you?

OpenStudy (anonymous):

not yet, im still catching up though

myininaya (myininaya):

\[a_3=\frac{-a_2}{3} \\a_4=\frac{-a_1}{12}+\frac{a_2}{6} \\ a_5=\frac{-a_2}{6}+\frac{a_1}{15} \\ y(x)=\sum_{n=0}^{\infty} a_n(x-1)^n \\ y(x)=a_0+a_1(x-1)+a_2(x-1)^2+a_3(x-1)^3+a_4(x-1)^4+a_5(x-1)^5+ \cdots \\ =a_0+a_1(x-1)+a_2(x-1)-\frac{a_2}{3}(x-1)^3+(\frac{-a_1}{12}+\frac{a_2}{6})(x-1)^4+ \\ +(\frac{-a_2}{6}+\frac{a_1}{15})(x-1)^5+ \cdots \]

myininaya (myininaya):

http://tutorial.math.lamar.edu/Classes/DE/SeriesSolutions.aspx I think we should just write this solution like they did in example 3

myininaya (myininaya):

so you would "combine like terms" get your a_0's together your a_1's and your a_2's

myininaya (myininaya):

I don't see a more elegant way

myininaya (myininaya):

but I could just be burnt out on this problem

OpenStudy (anonymous):

I dont blame you, i appreciate your help though!

myininaya (myininaya):

did you have any questions on this question before I leave?

OpenStudy (anonymous):

None thus far, I'm understanding everything and getting the same results when i work them out.

myininaya (myininaya):

like did you see that example 3 on paul's note our solution is going to look at crappy as that which is odd I wanted to find a pattern but I don't see how to do that here

OpenStudy (anonymous):

thank you for your help! :)

OpenStudy (anonymous):

Yeah i was just going over that, that matches with the solution in the back of the book so i think thats all we can do

myininaya (myininaya):

really?

myininaya (myininaya):

that's awesome and kind of a relief giving all the work that this problem took

OpenStudy (anonymous):

yep, they didn't get a general formula for the series either

myininaya (myininaya):

anyways goodnight

OpenStudy (anonymous):

goodnight and thank you again! I'd give you many more medals if I could lol

myininaya (myininaya):

it's cool :)